OpenStudy (anonymous):

a 200 N force acts on a 50kg ice skater. ignoring any friction, find the ice skater's acceleration. how do i set-up and solve this problem?

OpenStudy (anonymous):

Well, start from the Newton's law which say that the force acting on an object is equal to the mass of the object multiplied by the acceleration of the object. So, \[\ F=ma\] Also note that Force is a vector so you need to specify the direction of the force.

OpenStudy (y2o2):

F = ma 200 = 50a a = 200/50 = 4 (m/s²)